# Edit Transfer: Learning Image Editing via Vision In-Context Relations

### v1

@@ 2 Related Work
Vision In-context learning. In-context learning—the process of inferring patterns from provided input-output pairs—has been widely adopted in LLMs  [16]. In the visual domain, in-context learning is initially explored for understanding tasks [26, 27, 28, 29]. For instance, [29] demonstrates that in-context learning is effective for downstream representation tasks, such as segmentation. In the field of visual generation, leveraging DiT-based text-to-image (T2I) models, IC-LoRA [18] reveals that these models naturally exhibit in-context generation abilities. Building on these advances, we extend the visual in-context learning paradigm to Edit Transfer in this work, thereby developing a visual relation in-context learning approach.
@@ 3.2 Visual Relation In-Context Learning
P2P [20] RF-Solver-Edit [23] Mimic-Brush [13]
CLIP-T ($\uparrow$) 20.22 19.83 - 22.58
CLIP-I ($\uparrow$) - - 0.804 0.810
PickScore ($\uparrow$) 20.60 20.96 20.34 21.50
@@ 4.3 Comparisons with Baselines
  • TIE methods: 1) P2P [20] is a classical text-based editing method; 2) RF-Solver-Edit [23] is a state-of-the-art editing method that incorporates MasaCtrl [22]’s principles into the DiT architecture, delivering enhanced editing capabilities.
  • RIE method: MimicBrush [13] enables users to edit image regions by drawing on in-the-wild references without needing an exact source match.
@@ 4.4 Ablation Study
Figure 11: Limitations. Our method struggles with low-level properties transfer e.g. color.
@@ 4.5 Discussion
Limitations. While our proposed framework excels in spatial and compositional editing tasks, it struggles with low-level attribute transfer. As shown in Fig. 11, the model struggles to recolor a new T-shirt based on the editing transformation demonstrated in the source–target pair.

### v2

@@ 2 Related Work
Vision In-context learning. In-context learning—the process of inferring patterns from provided input-output pairs—has been widely adopted in LLMs  [16]. In the visual domain, in-context learning is initially explored for understanding tasks [26, 27, 28, 29]. For instance, [29] demonstrates that in-context learning is effective for downstream representation tasks, such as segmentation. In the field of visual generation, leveraging DiT-based text-to-image (T2I) models, IC-LoRA [18] reveals that these models naturally exhibit in-context generation abilities. It focuses on generating a set of sub-images with a consistent style or synthesizing images conditioned on a reference image. In contrast, our work aims to address the proposed edit transfer task by extending the visual in-context learning paradigm, enabling the model to infer and apply transformations demonstrated in an editing exemplar to query images.
@@ 3.2 Visual Relation In-Context Learning
P2P [20] RFSE [23] SF [24] MB [13] PX [14]
CLIP-T $\uparrow$ 20.22 19.83 18.67 - - 22.58
CLIP-I $\uparrow$ - - - 0.804 0.761 0.810
PickScore $\uparrow$ 20.60 20.96 20.91 20.34 20.82 21.50
@@ 4.3 Comparisons with Baselines
  • TIE methods: 1) P2P [20] is a classical text-based editing method; 2) RF-Solver-Edit (RFSE) [23] and StableFlow (SF) [24] propose new inversion methods for flow matching and adapt MasaCtrl [22]’s principles into the DiT architecture, delivering enhanced editing capabilities.
  • RIE method: MimicBrush (MB) [13] and PIXELS (PX) [14] enable users to edit image regions by drawing on in-the-wild references without needing an exact source match.
@@ 4.4 Ablation Study
Figure 11: Applications to appearance and style transfer. Our proposed method can also be applied to other editing types such as appearance and style transfer.
@@ 4.5 Discussion
Application to other editing types. While our focus is on non-rigid editing tasks, the proposed EditTransfer can also be extended to other editing types such as appearance and style transfer, as demonstrated in Fig. 11.